### v4

Abstract
Not every directed acyclic graph (DAG) whose underlying undirected graph is planar admits an upward planar drawing. We are interested in pushing the notion of upward drawings beyond planarity by considering upward $k$-planar drawings of DAGs in which the edges are monotonically increasing in a common direction and every edge is crossed at most $k$ times for some integer $k\geq 1$. We show that the number of crossings per edge in a monotone drawing is in general unbounded for the class of bipartite outerplanar, cubic, or bounded pathwidth DAGs. However, it is at most two for outerpaths and it is at most quadratic in the bandwidth in general. From the computational point of view, we prove that testing upward $k$-planarity is NP-complete already for $k=1$ and even for restricted instances for which upward planarity testing is polynomial. On the positive side, we can decide in linear time whether a single-source DAG admits an upward $1$-planar drawing in which all vertices are incident to the outer face.
keywords:
Upward drawings, beyond planarity, upward $k$-planarity, upward outer-1-planarity
1 Introduction
Graph drawing “beyond planarity” studies the combinatorial and algorithmic questions related to representations of graphs where edges can cross but some crossing configurations are forbidden. Depending on the forbidden crossing configuration, different beyond-planar types of drawings can be defined including, for example, RAC, $k$-planar, fan planar, and quasi planar drawings (see Didimo, Liotta, and Montecchiani, 2019; Hong and Tokuyama, 2020; Kobourov, Liotta, and Montecchiani, 2017, for surveys and books).
While most of the literature about beyond planar graph drawing has focused on undirected graphs (one of the few exceptions being the paper of Angelini, Cittadini, Didimo, Frati, Di Battista, Kaufmann, and Symvonis (2011), which studies RAC upward drawings), we study upward $k$-planar drawings of acyclic digraphs (DAGs), i.e., drawings of DAGs where the edges monotonically increase in $y$-direction and each edge can be crossed at most $k$ times. The minimum $k$ such that a DAG admits an upward $k$-planar drawing is called its upward local crossing number. We focus on values of $k=1,2$ and investigate both combinatorial properties and complexity questions. Our research is motivated by the observation that well-known DAGs that are not upward planar, i.e., not upward 00-planar, do admit a drawing where every edge is crossed at most a constant number of times; see, e.g., Fig. 1.
Figure 1: A graph that is not upward planar but admits an upward 1-planar drawing.
Our contribution.
  • A graph is an outerpath if it has a planar drawing in which each vertex is incident to the outer face and the internal faces induce a path in the dual graph. Papakostas (1994) observed that there is a directed acyclic 8-vertex outerpath that is not upward-planar (see Fig. 3(a)). We strengthen this observation by showing that there exists a directed acyclic fan (that is, a very specific outerpath) that has no upward-planar drawing (Proposition 1). On the other hand, we show that every directed acyclic outerpath is upward 2-planar (Theorem 4) and that the upward local crossing number is quadratic in the bandwidth (Theorem 3). However, the upward local crossing number of bipartite outerplanar DAGs (Theorem 1), bipartite DAGs with bounded pathwidth (Corollary 1), and cubic DAGs (Proposition 2) is in general unbounded.
  • We show that upward $1$-planarity testing is NP-complete, even for graph families where upward planarity testing can be solved in polynomial time. These include: single-source single-sink series-parallel DAGs with a fixed rotation system; single-source two-sink series-parallel DAGs where the rotation system is not fixed; and single-source single-sink DAGs without fixed rotation system that can be obtained from a $K_{4}$ by replacing the edges with series-parallel DAGs (Theorem 6).
  • Finally, following a common trend in the study of beyond planar graph representations, we consider the outer model, in which all vertices are required to lie on a common face while maintaining the original requirements (Didimo, Liotta, and Montecchiani, 2019; Hong and Tokuyama, 2020; Kobourov, Liotta, and Montecchiani, 2017). We prove that testing whether a single-source DAG admits an upward outer-1-planar drawing can be done in linear time (Theorem 7).
Related Work.
A drawing of a graph is monotone if all edges are drawn monotone with respect to some direction, e.g., a drawing is y-monotone or upward, if each edge intersects each horizontal line at most once. The corresponding crossing number is introduced and studied by Valtr (2005) and Fulek, Pelsmajer, Schaefer, and Štefankovič (2013). Schaefer (2024) mentions the upward crossing number and the local crossing number but not their combination. Schaefer (2024, p. 64) also showed that a drawing with the minimum number of crossings per edge can require incident edges that cross. The edges of the provided 4-planar example graph can be oriented such that the resulting directed graph admits an upward 4-planar drawing. Thus, also an upward drawing that achieves the minimum local crossing number can require incident edges that cross. Also, the so-called strong Hanani–Tutte theorem carries over to directed graphs: Fulek et al. (2013, Theorem 3.1) showed that every undirected graph that has a monotone drawing where any pair of independent edges crosses an even number of times also has a planar monotone drawing with the same vertex positions. This implies that in any upward drawing of a graph that is not upward-planar there must be a pair of independent edges that crosses an odd number of times.
Upward drawings of directed acyclic graphs have been studied in the context of (upward) book embeddings. In that model the vertices are drawn on a vertical line (a spine) following a topological order of the graph, while all edges are pointing upwards. To reduce the edge crossings, edges are partitioned into the fewest number of crossing-free subsets (pages). Studying upward book embeddings is a popular topic, which is usually centered around determining the smallest number of pages for various graph classes (Heath, Pemmaraju, and Trenk, 1999; Frati, Fulek, and Ruiz-Vargas, 2011, 2013; Nöllenburg and Pupyrev, 2023; Jungeblut, Merker, and Ueckerdt, 2023) or deciding whether a graph admits an upward drawing with a given number of pages (Binucci, Da Lozzo, Di Giacomo, Didimo, Mchedlidze, and Patrignani, 2023; Bhore, Da Lozzo, Montecchiani, and Nöllenburg, 2023; Bekos, Da Lozzo, Frati, Gronemann, Mchedlidze, and Raftopoulou, 2023). Our model is related to topological book embeddings (Mchedlidze and Symvonis, 2009; Giordano, Liotta, Mchedlidze, Symvonis, and Whitesides, 2015), which are a relaxed version of 2-page book embeddings in which edges are allowed to cross the spine. While papers about topological book embeddings insist on planar drawings and minimize the number of spine crossings, we do allow crossings and want to bound the maximum number of crossings per edge (ignoring the spine).
2 Preliminaries
A drawing $\Gamma$ of a graph $G$ maps the vertices of $G$ to distinct points in the plane and the edges of $G$ to open Jordan curves connecting their respective endpoints but not containing any other vertex point. A crossing between two edges is a common point of their curves, other than a common end point. A drawing is simple if (a) there are no self-crossings (the Jordan curves are simple), (b) no two edge curves share more than one point, and (c) no three edge curves share a common internal point.
For a vertex $v$ of $G$ and a drawing $\Gamma$ of $G$, let $x_{\Gamma}(v)$ and $y_{\Gamma}(v)$ denote the x- and y-coordinates of $v$ in $\Gamma$, respectively; when $\Gamma$ is clear from the context, we may omit it and simply use the notation $x(v)$ and $y(v)$. A face of $\Gamma$ is a region of the plane delimited by maximal uncrossed segments of the edges of $G$. The unique unbounded face of $\Gamma$ is its outer face, the other faces are its internal faces. An outer edge is one incident to the outer face; all other edges are inner edges. The rotation of a vertex $v$ in $\Gamma$ is the counterclockwise cyclic order of the edges incident to $v$. The rotation system of $\Gamma$ is the set of rotations of its vertices.
The drawing $\Gamma$ is planar if no two of its edges cross; it is $k$-planar if each edge is crossed at most $k$ times. A graph is ($k$-)planar if it admits a ($k$-)planar drawing; it is outer ($k$-)planar if it admits a ($k$-)planar drawing where all vertices are incident to the outer face. A planar embedding $\mathcal{E}$ of a planar graph $G$ is an equivalence class of planar drawings of $G$, namely those that have the same set of faces. Each face can be described as a sequence of edges and vertices of $G$ which bound the corresponding region in the plane; each such sequence is a face of $G$ in the embedding $\mathcal{E}$. A planar embedding $\mathcal{E}$ of a connected graph can also be described by specifying the rotation system and the outer face associated with any drawing of $\mathcal{E}$.
Let $\Gamma$ be a non-planar drawing of a graph $G$; the planarization of $\Gamma$ is the planar drawing $\Gamma^{\prime}$ of the planarized graph $G^{\prime}$ obtained by replacing each crossing of $\Gamma$ with a dummy vertex. If $\Gamma$ is 1-planar, the planarization can be obtained as follows. Let $uv$ and $wz$ be any two edges that cross in $\Gamma$; they are replaced in $\Gamma^{\prime}$ by the edges $ux$, $xv$, $wx$ and $xz$, where $x$ is the dummy vertex. Two non-planar drawings of a graph $G$ have the same embedding if their planarizations have the same planar embedding. An embedding $\mathcal{E}$ of $G$ can also be described by specifying the planarized graph $G^{\prime}$ and one of its planar embeddings. A planar graph with a given planar embedding is also called plane graph. An outerplane graph is a plane graph whose vertices are all incident to the outer face. A fan is a maximal outerpath that contains a vertex $c$ that is adjacent to all other vertices; we call $c$ the central vertex of the fan. A $2$-tree is a planar graph that can be reduced to an edge by iteratively removing a degree-two vertex that closes a $3$-cycle. A series-parallel graph is a graph that can be augmented to a $2$-tree by adding edges (and no vertices).
A (simple, finite) directed graph (digraph for short) $G$ consists of a finite set $V(G)$ of vertices and a finite set $E(G)\subseteq\{(u,v)\mid u,v\in V(G),u\neq v\}$ of ordered pairs of vertices, which are called edges. A source (resp. sink) of $G$ is a vertex with no incoming (resp. no outgoing) edges. A single-source graph is a digraph with a single source and, possibly, multiple sinks. A digraph $G$ is an st-graph if: (i) it is acyclic and (ii) it has a single source $s$ and a single sink $t$. An st-graph is a planar st-graph if it admits a planar embedding with $s$ and $t$ on the outer face. We say that a drawing of a digraph $G$ is upward if every (directed) edge $(u,v)$ of $G$ is mapped to a $y$-monotone Jordan arc with $y(u)<y(v)$. Clearly, a digraph admits an upward drawing only if it does not contain a directed cycle. Therefore, we assume for the rest of the paper that the input graph is a DAG, a directed acyclic graph. Such a graph has a linear extension, i.e., a vertex order $v_{1},\dots,v_{n}$ such that, for any directed edge $(v_{i},v_{j})$, we have $i<j$. We say that a DAG is planar, outerplanar, or series-parallel if its underlying undirected graph is planar, outerplanar, or series-parallel, respectively.
Let $\Gamma$ be an upward drawing of a DAG $G$. By the upwardness, the rotation system of $\Gamma$ is such that for every vertex $v$ of $\Gamma$ the rotation of $v$ has only one maximal subsequence of outgoing (incoming) edges. We call such a rotation system a bimodal rotation system. An upward embedding of a DAG $G$ is an embedding of $G$ arising from an upward drawing; it naturally has a bimodal rotation system. The minimum $k$ such that a digraph $G$ admits an upward $k$-planar drawing is called its upward local crossing number and denoted by $\operatorname{lcr^{\uparrow}}(G)$.
For any positive integer $k$, we use $[k]$ as shorthand for $\{1,2,\dots,k\}$. A path-decomposition of a graph $G$ is a sequence $P=\left<X_{1},\dots,X_{\ell}\right>$ of subsets of $V(G)$, called buckets, such that (1) for each edge $e$ of $G$ there is a bucket that contains both end vertices of $e$, and (2) for every vertex $v$ of $G$, the set of buckets that contain $v$ form a contiguous subsequence of $P$. The width of a path-decomposition is one less than the size of the largest bucket. The pathwidth of the graph $G$ is the width of a path decomposition of $G$ with the smallest width.
3 Lower Bounds
We start with a negative result that shows that even very special directed acyclic outerpaths may not admit upward-planar drawings, thus strengthening Papakostas’ observation (1994).
Proposition 1.
Not every directed acyclic fan is upward-planar.
Proof.
Consider the 7-vertex fan $F$ depicted in Fig. 2(a). Suppose for a contradiction that $F$ is upward planar, that is, $F$ admits an upward planar drawing $\Gamma$. Let $c$ be the central vertex of $F$. We assume that $c$ is placed at the origin. We say that a triangle of $F$ is positive (negative, respectively) if the corresponding region of the plane in $\Gamma$ contains the point $(\varepsilon,0)$ ($(-\varepsilon,0)$, respectively) for a sufficiently small value $\varepsilon>0$. The triangles that have one vertex below $c$ and one vertex above $c$ (namely $t_{1}=\triangle cv_{1}v_{2}$, $t_{3}=\triangle cv_{3}v_{4}$, and $t_{5}=\triangle cv_{5}v_{6}$) are either positive or negative.
(a) the fan $F$
(b) case: $t_{5}$ is contained in $t_{1}$
(c) a 1-planar upward drawing of $F$
Figure 2: A directed acyclic fan $F$ that does not admit an upward planar drawing.
If both $t_{1}$ and $t_{5}$ are positive, then one must contain the other in $\Gamma$, say, $t_{1}$ contains $t_{5}$; see Fig. 2(b). But then vertices $v_{3}$ and $v_{4}$ must also lie inside $t_{1}$. If both lie inside $t_{5}$, then the edge $(v_{3},v_{2})$ intersects an edge of $t_{5}$. If one of them lies inside $t_{5}$ and one does not, then the edge $(v_{4},v_{3})$ intersects an edge of $t_{5}$. So both must lie outside $t_{5}$. But $v_{4}$ lies on one hand above $v_{5}$ and on the other hand below $c$ and, thus, below $v_{6}$. So the edge $(v_{4},c)$ intersects the edge $(v_{5},v_{6})$. (If $t_{5}$ contains $t_{1}$, the edge $(c,v_{3})$ intersects the edge $(v_{1},v_{2})$.)
By symmetry, not both $t_{1}$ and $t_{5}$ can be negative, so exactly one of $t_{1}$ and $t_{5}$ must be negative, say, $t_{1}$; see Fig. 2(c). Now first assume that $t_{3}$ is positive. Due to edge $(v_{3},v_{2})$, vertex $v_{3}$ must be outside $t_{5}$, so $t_{3}$ cannot be inside $t_{5}$. On the other hand, $t_{3}$ cannot contain $t_{5}$ because $v_{4}$ is above $v_{5}$. Hence $t_{3}$ intersects $t_{5}$. Finally, assume that $t_{3}$ is negative. Due to edge $(v_{5},v_{4})$, vertex $v_{4}$ must be outside $t_{1}$, so $t_{3}$ cannot be inside $t_{1}$. On the other hand, $t_{3}$ cannot contain $t_{1}$ because $v_{3}$ is below $v_{2}$. Hence $t_{3}$ intersects $t_{1}$. ∎
By iteratively adding paths on every outer edge of an outerplanar but not upward-planar DAG, we can construct outerplanar DAGs with an unbounded upward local crossing number.
Theorem 1.
For each $\ell\geq 0$, there is a bipartite outerplanar DAG $G_{\ell}$ with $n_{\ell}=8\cdot 3^{\ell}$ vertices, maximum degree $\Delta_{\ell}=2\ell+3$, and upward local crossing number greater than $\ell/6$, which is in $\Omega(\log n_{\ell})$ and $\Omega(\Delta_{\ell})$.
Proof.
The bipartite graph $G_{0}$ in Fig. 3(a) is not upward planar (Papakostas, 1994). For $\ell\geq 1$, we construct $G_{\ell}$ from $G_{\ell-1}$ by adding a $3$-edge path on every outer edge of the graph. Fig. 3(b) shows $G_{2}$. The maximum degree of $G_{\ell}$ is $\Delta_{\ell}=2\ell+3$. The number of vertices is $n_{\ell}=8+\sum_{i=1}^{\ell}8\cdot 3^{i-1}\cdot 2=8\cdot 3^{\ell}$.
(a) $G_{0}$
(b) $G_{2}$
Figure 3: There is a family $(G_{\ell})_{\ell\geq 0}$ of bipartite outerplanar graphs such that $G_{\ell}$ has $n_{\ell}$ vertices, maximum degree $\Delta_{\ell}$, and upward local crossing number in $\Omega(\Delta_{\ell})\cap\Omega(\log n_{\ell})$.
(a) $e^{\prime}$ inner edge of $G_{0}$
(b) $e^{\prime}$ outer edge of $G_{0}$
Figure 4: If $e$ crosses $e^{\prime}$ an odd number of times then there is a cycle $C$ of length at most $6$ that is crossed by at least $\ell+1$ edge-disjoint paths, namely the edge $e$ and the $\ell$ paths added on top of $e$.
Consider now an upward $k$-planar drawing $\Gamma$ of $G_{\ell}$ for some $k\leq\ell$. Since $G_{0}$ is not upward planar, there must be a pair of independent edges of $G_{0}$ that crosses an odd number of times in $\Gamma$. Observe that $G_{0}$ has no upward planar drawing in which only the two inner edges cross an odd number of times, for otherwise the two cycles $\langle 1,2,6,5\rangle$ and $\langle 3,4,8,7\rangle$ would intersect an odd number of times, which is impossible. Thus, in $\Gamma$ there must be an outer edge $e$ of $G_{0}$ that is crossed by an independent edge $e^{\prime}$ of $G_{0}$ an odd number of times. We choose $e^{\prime}$ to be an outer edge of $G_{0}$, if possible.
We now determine a cycle $C$ of $G_{\ell}$ that is crossed by $e$ an odd number of times and does also not contain any end vertex of $e$. If $e^{\prime}$ is an inner edge, then we take the outer path $P$ of $G_{0}$ that connects the ends of $e^{\prime}$ and does not contain $e$; this is not crossed by $e$ due to our choice of $e^{\prime}$. Moreover, since $e^{\prime}$ and $e$ are independent, it follows that $P$ and $e$ do not share a vertex. Let $C$ be the concatenation of $P$ and $e^{\prime}$. Note that the length of $C$ is at most six; see Fig. 4(a).
If $e^{\prime}$ is an outer edge of $G_{0}$, we do the following: We start with the path $P$ of length three that was added for $e^{\prime}$. Since $e$ is an edge of $G_{0}$ and $e$ and $e^{\prime}$ are independent, it follows that $P$ and $e$ do not share a vertex. If $P$ contains an edge that is crossed an odd number of times by $e$ then we continue with such an edge instead of $e^{\prime}$. More precisely, let $e_{1}=e^{\prime}$ and initialize $i=1$. Let $P_{i}$ be the path of length three that was added for $e_{i}$. While $P_{i}$ contains an edge that is crossed an odd number of times by $e$, let $e_{i+1}$ be such an edge, let $P_{i+1}$ be the path of length three that was added for $e_{i}$, and increase $i$ by one. Since $e$ is crossed at most $k$ times, this process stops at some $i\leq k$. Let $C$ be the cycle that is composed of $P_{i}$ and $e_{i}$. See Fig. 4(b). In this case $C$ has length four. Moreover $C$ shares at most the end vertices of $e^{\prime}$ with $G_{0}$. Thus, since $e$ is an edge of $G_{0}$ and $e$ and $e^{\prime}$ are independent, it follows that $C$ and $e$ do not share a vertex.
Cycle $C$ might cross itself. However, it divides the plane into cells. Since $e$ crosses $C$ an odd number of times, it follows that the end vertices of $e$ must be in different cells of the plane. This means that not only $e$ but also the $\ell$ edge-disjoint paths that were added on top of $e$ have to cross $C$. Observe that none of these paths contains a vertex of $C$. But $C$ contains at most six edges, each of which can be crossed at most $k$ times. This is impossible if $\ell\geq 6k$. Hence, if there is an upward $k$-planar drawing then $\ell<6k$, which means that $k>\ell/6$. ∎
We now show that if we expand the graph class beyond outerplanar graphs, then we get a lower bound on the upward local crossing number that is even linear in the number of vertices. The graphs in our construction have pathwidth 2, as opposed to the graphs in Theorem 1 whose pathwidth is logarithmic. Observe that a caterpillar, i.e., a tree that can be reduced to a path by removing all degree-1 vertices, has pathwidth 1, and that the pathwidth can increase by at most 1 if we add a vertex with some incident edges or subdivide some edges.
Theorem 2.
For every $k\geq 1$, there exists a (planar) DAG with $\Theta(k)$ vertices, maximum degree in $\Theta(k)$, and pathwidth 2 that does not admit an upward $k$-planar drawing.
Proof.
Let $G_{k}$ be the graph consisting of the four vertices $a$, $b_{1}$, $b_{2}$, and $c$ and the following set of edges and degree-2 vertices (see also Fig. 5):
  • edges $(a,b_{1})$ and $(a,b_{2})$;
  • for $i\in[2]$ and $j\in[3k+1]$, a through-vertex at $b_{i}$, i.e., a vertex $d_{i}^{(j)}$ and edges $(b_{i},d_{i}^{(j)})$ and $(d_{i}^{(j)},c)$;
  • for $j\in[6k+1]$, a source below $a$, i.e., a vertex $s^{(j)}$ and edges $(s^{(j)},a)$ and $(s^{(j)},c)$;
  • for $i\in[2]$ and $j\in[4k+1]$, a sink above $b_{i}$, i.e., a vertex $t_{i}^{(j)}$ and edges $(b_{i},t_{i}^{(j)})$ and $(c,t_{i}^{(j)})$.
Clearly, $G_{k}$ has $\mathcal{O}(k)$ vertices, and pathwidth 2, since $G-c$ is a caterpillar and has pathwidth 1.
Figure 5: A graph of pathwidth 2 (drawn upward) that does not have an upward $k$-planar drawing. (a) We only show three of the $\Theta(k)$ vertices of each group. (b) Cycles $C_{a}$ and $C_{b}$.
Assume that there was an upward $k$-planar drawing $\Gamma$ of $G_{k}$. Up to renaming, we may assume that $y(b_{2})\leq y(b_{1})$. Delete all but one of the through-vertices at $b_{1}$ from the drawing; in what follows, we write $d_{1}$ for the one that we keep (it does not matter which one).
Among the $13k+$ 1 through-vertices $d_{2}^{(j)}$ at $b_{2}$, there exists at least one for which the path $\langle b_{2},d_{2}^{(j)},c\rangle$ crosses none of the three edges in the path $\langle a,b_{1},d_{1},c\rangle$, for otherwise there would be an edge with more than $k$ crossings. Delete all other through-vertices at $b_{2}$; in what follows we write $d_{2}$ for the one that we keep. Let $a^{\prime}$ be the topmost intersection point of $(a,b_{1})$ and $(a,b_{2})$ (possibly $a^{\prime}=a$). Since $y(a)\leq y(a^{\prime})<y(b_{2})\leq y(b_{1})$ the curve $C_{b}$ formed by the two directed paths $\langle a^{\prime},b_{i},d_{i},c\rangle$ (for $i\in[2]$) is drawn without crossing in $\Gamma$.
Curve $C_{b}$ uses six edges, therefore among the $16k+$ 1 sources below $a$, there exists one, call it $s$, for which edge $(s,c)$ crosses no edge of $C_{b}$. Since $y(s)<y(a)$, vertex $s$ is outside $C_{b}$, and so the entire edge $(s,c)$ is outside $C_{b}$, except at the endpoint $c$. In particular, among the three edges $(d_{1},c)$, $(d_{2},c)$, and $(s,c)$ that are incoming at $c$, edge $(s,c)$ is either leftmost or rightmost (but cannot be the middle one). We assume here that $(s,c)$ is rightmost, the other case is symmetric. Write $\{p,q\}=\{1,2\}$ such that the left-to-right order of incoming edges at $c$ is $(d_{p},c)$, $(d_{q},c)$, $(s,c)$. In Fig. 5, we have $p=1$ and $q=2$.
Edge $(s,a)$ is also outside $C_{b}$, except perhaps at endpoint $a$, since it uses smaller $y$-coordinates. Let $s^{\prime}$ be the topmost intersection point of $(s,a)$ and $(s,c)$. Then there are no crossings in the curve $C_{a}$ formed by the directed paths $\langle s^{\prime},a,b_{p},d_{p},c\rangle$ and $\langle s^{\prime},c\rangle$. By our choice of $p$ and $q$, vertex $d_{q}$ is inside $C_{a}$, and so is the entire path $\langle a^{\prime},b_{q},d_{q},c\rangle$, except at the ends since it is part of $C_{b}$. In particular, $b_{q}$ is inside $C_{a}$, whereas, for $j\in[4k+1]$, $t_{q}^{(j)}$ is outside $C_{a}$ due to $y(c)<y(t_{q}^{(j)})$. It follows that one of the four edges $(a,b_{p})$, $(b_{p},d_{p})$, $(d_{p},c)$ and $(s,c)$ must be crossed at least $k+1$ times by edges from $b_{q}$ to the sinks above it. Thus, the drawing was not $k$-planar, a contradiction. ∎
The graphs that we constructed in the proof of Theorem 2 are not bipartite, but one can make them bipartite by subdividing all edges once. This at best cuts the local crossing number in half, increases the pathwidth by at most 1, and yields the following result.
Corollary 1.
There is a family of bipartite (planar) DAGs of constant pathwidth whose upward local crossing number is linear in the number of vertices.
So far we needed graphs of unbounded maximum degree in order to enforce unbounded upward local crossing number. We now show that, intrinsically, this is not necessary.
Proposition 2.
There are cubic DAGs whose upward local crossing number is at least linear in the number of vertices.
Proof.
The crossing number of a random cubic graph with $n$ vertices is expected to be at least $cn^{2}$ for some absolute constant $c>0$ (Dujmović, Kawarabayashi, Mohar, and Wood, 2008), and thus there exist graphs yielding this bound. By the pigeon-hole principle, such a graph contains an edge with $\Omega(n)$ crossings among its $\Theta(n)$ edges. Impose arbitrary acyclic edge directions. ∎
4 Upper Bounds
The bandwidth $\operatorname{bw}(G)$ of an undirected graph $G$ is the smallest positive integer $k$ such that there is a labeling of the vertices by distinct numbers $1,\dots,n$ for which the labels of every pair of adjacent vertices differ by at most $k$.
Theorem 3.
The upward local crossing number of a DAG $G$ with maximum degree $\Delta$ is at most $\Delta\cdot(\operatorname{bw}(G)-2)\leq 2\operatorname{bw}(G)(\operatorname{bw}(G)-2)$, so it is in $\mathcal{O}(\Delta\cdot\operatorname{bw}(G))\subseteq{\mathcal{O}}(\operatorname{bw}(G)^{2})$.
Proof.
Observe that the maximum degree $\Delta$ of a graph $G$ is bounded in terms of the bandwidth of $G$; namely, $\Delta\leq 2\operatorname{bw}(G)$. Consider a linear extension of $G$. For every vertex $v$ of $G$, let $y(v)$ be its index in the extension. Now consider a labeling of $G$ corresponding to the bandwidth. For every vertex $v$ of $G$, let $x(v)$ be its label.
Construct a drawing of $G$ by first placing every vertex $v$ at the point $(x(v),y(v))$ and by then perturbing vertices slightly so that the points are in general position. Adjacent vertices are connected via straight-line segments.
It is easy to see that the drawing is upward since it is consistent with the linear extension. Consider an arbitrary edge $(u,v)$ and assume, without loss of generality, that $x(u)<x(v)$. Let $\ell=x(v)-x(u)$. The length of any edge in x-direction is bounded by $\operatorname{bw}(G)$; hence $\ell\leq\operatorname{bw}(G)$. Edge $(u,v)$ may be crossed (a) by edges that have at least one incident vertex $w$ with $x(u)<x(w)<x(v)$, or (b) by edges between two vertices $w$ and $w^{\prime}$ with $x(w)<x(u)$ and $x(v)<x(w^{\prime})$. Thus, there are at most $\ell-1$ and $\operatorname{bw}(G)-\ell-1$ possible choices for $w$ in the two scenarios, respectively. Since each vertex is incident to at most $\Delta$ edges, $(u,v)$ can be crossed at most $\Delta(\operatorname{bw}(G)-2)$ times. ∎
For some graph classes, sublinear bounds on the bandwidth are known (see, e.g., Böttcher, Pruessmann, Taraz, and Würfl, 2010; Fomin and Golovach, 2003; Wood, 2006). This gives upper bounds on the upward local crossing number of many graph classes. We list only a few:
Corollary 2.
The following classes of DAGs have sublinear upward local crossing number:
  • Square $k\times k$ grids have bandwidth $\Theta(k)$ and $\Delta=4$, hence their upward local crossing number is in $\mathcal{O}(k)=\mathcal{O}(\sqrt{n})$.
  • Planar graphs of maximum degree $\Delta$ have bandwidth ${\mathcal{O}}({n}/{\log_{\Delta}n})$ (Böttcher et al., 2010), hence their upward local crossing number is in $\mathcal{O}({n\cdot\Delta}/{\log_{\Delta}n})$. Observe that ${n\cdot\Delta}/{\log_{\Delta}n}\in o(n)$ if $\Delta\in o({\log n}/{\log\log n})$.
We now complement the negative result in Proposition 1 by showing that every directed acyclic outerpath allows an upward 2-planar drawing. We start with a technical lemma on fans.
(c) a directed fan $G$
(d) upward drawing of $G$ with at most two crossings per edge
Figure 6: Constructing upward $2$-planar drawings of fans according to Lemma 1.
Lemma 1.
Let $c$ be the central vertex of a directed acyclic fan $G$, and let $P=\langle v_{1},\dots,v_{n-1}\rangle$ be the path of the remaining vertices in $G$. Let $P_{1},\dots,P_{k}$ be an ordered partition of $P$ into maximal subpaths such that, for every $i\in[k]$, the edges between $P_{i}$ and $c$ either are all directed towards $c$ or are all directed away from $c$. Then there is an upward 2-planar drawing of $G$ with the following properties:
  1. 1.
    no edge incident to $c$ is crossed;
  2. 2.
    the central vertex $c$ and $v_{n-1}$ have x-coordinate $n-1$, and the x-coordinates of $v_{1},\dots,v_{n-2}$ are pairwise distinct values within $\{1,\dots,n-2\}$;
  3. 3.
    for all edges all $x$-coordinates of the curves are at most $n-1$; all edges incident to $c$ and all edges of the subpaths $P_{1},\dots,P_{k}$ are in the vertical strip between $1$ and $n-1$;
  4. 4.
    if $P_{1}$ is a directed path, then the edge between $P_{1}$ and $P_{2}$ is crossed at most once.
Proof.
We place $c$ at $(n-1,0)$; then we place $v_{1},v_{2},\dots,v_{n-1}$ above or below $c$ depending on the direction of the edges that connect them to $c$; see Fig. 6 for an example.
For each $\ell\in[n-2]$, we keep the invariant that, when we place $v_{\ell}$, the leftward ray, that is, the one in direction ($1\choose 0$ from $v_{\ell}$, reaches the outer face of the current drawing after crossing at most one other edge, and that this edge is currently crossed by at most one edge.
In order to choose appropriate y-coordinates, we maintain two values $y_{\min}$ and $y_{\max}$ indicating the minimum and maximum y-coordinate of any so far drawn vertex. Consider now a subpath $P^{\prime}\in\{P_{1},\dots,P_{k}\}$. Let $v_{h}$ be the first and let $v_{j-1}$ be the last vertex of $P^{\prime}$, i.e., $P^{\prime}=\langle v_{h},v_{h+1},\dots,v_{j-1}\rangle$. We describe in detail the case that the edge from $v_{h},\dots,v_{j-1}$ to $c$ are directed towards $c$ that is, $v_{h}$ must lie below $c$. The other case is symmetric. We place $v_{h}$ at x-coordinate $h$ and with a y-coordinate sufficiently below $y_{\min}$. If $h=j-1$ we are done.
We now consider the case that $j=n$ or $(v_{j-1},v_{j-2})\in E(G)$; see, for example, path $P_{5}$ in Fig. 6. In this case, we place $v_{h+1},\dots,v_{j-1}$ using x-coordinates $h+1,\dots,j-1$, going up and down as needed but remaining below the x-axis. The edges are drawn such that all vertices of $P^{\prime}$ remain on the outer face of the drawing. I.e., if we use straight-line edges, then, for $i\in[n-2]$, the slope of $v_{i}v_{i+1}$ must be less than the slope of $v_{i}c$. Since we go towards $c$, we can draw $P^{\prime}$ and the edges that connect $v_{1},v_{2},\dots,v_{n-1}$ to $c$ without any crossings.
If $j\neq n$ and $(v_{j-2},v_{j-1})\in E(G)$, then let $i\in\{h,\dots,j-1\}$ be the smallest index such that the subpath $\langle v_{i},v_{i+1},\dots,v_{j-1}\rangle$ is directed. See, for example, that last part of $P_{1}$ in Fig. 6. In that case, we place $v_{h+1},\dots,v_{i-1}$ at x-coordinates $h+1,\dots,i-1$, going slightly up and down as in the case described above. Let $y_{\min}$ be the smallest among the y-coordinates of all points placed so far.
Then we place $v_{i},v_{i+1},\dots,v_{j-1}$ in reverse order, i.e., at x-coordinates $j-1,j-2,\dots,i$. For the y-coordinates, we choose $y(v_{i})=y_{\min}-\gamma$ and $y(v_{j-1})=y_{\min}-\varepsilon$ for some (large) $\gamma>0$ and (small) $\varepsilon>0$ with the following properties: if $i>h$ then $v_{j-1}$ lies inside the triangle $\triangle v_{i-1}v_{i}c$ (pale yellow in Fig. 6) and if $i=h$ then $v_{j-1}$ lies inside the triangle $\triangle ov_{i}c$, where $o=(0,0)$. Draw $v_{i+1},\dots,v_{j-2}$ on the segment $\overline{v_{i}v_{j-1}}$. This fulfills the invariant: if $i>h$ then the vertex $v_{j-1}$ can reach the outer face via the edge $(v_{i},v_{i-1})$ which was not crossed so far. If $i=h$ then $v_{j-1}$ is on the outer face if $P^{\prime}=P_{1}$, otherwise $v_{j-1}$ can reach the current outer face by crossing the edge $(v_{h},v_{h-1})$. Observe that, by our invariant, $(v_{h},v_{h-1})$ might have crossed one edge in order to reach the outer face. While drawing $P^{\prime}$, we do not cross $(v_{h},v_{h-1})$ again. So the potential crossing with $(v_{j-1},v_{j})$ is at most the second crossing of $(v_{h},v_{h-1})$, and $(v_{h},v_{h-1})$ will not be crossed later.
Moreover, when we draw the next maximal subpath, we place $v_{j}$ at $(j,y_{\max}+1)$, i.e., in particular in the outer face of the current drawing. The edge from $v_{j-1}$ to $v_{j}$ must be directed towards $v_{j}$ since the orientation is acylic. Thus, we can draw the edge between $v_{j-1}$ and $v_{j}$ upward with at most one crossing, causing at most a second crossing on $(v_{h},v_{h-1})$. ∎
(a) a directed outerpath $G^{\prime}$
(b) upward drawing of $G^{\prime}$ with at most two crossings per edge
Figure 7: Example in- and output of our drawing algorithm (edge crossings due to Lemma 1 are highlighted in yellow; other edge crossings are highlighted in orange).
Now we describe our construction for general outerpaths; see Fig. 7.
Theorem 4.
Every directed acyclic outerpath admits an upward 2-planar drawing.
Proof.
Without loss of generality, we can assume that the given outerpath is maximal: if the outerpath has interior faces that are not triangles, we temporarily triangulate them using additional edges, which we direct such that they do not induce directed cycles and which we remove after drawing the maximal outerpath.
Let $G^{\prime}$ be such a graph; see Fig. 7. Let $c_{1},c_{2},\dots,c_{k}$ be the vertices of degree at least 4 in $G^{\prime}$ (marked red in Fig. 7). These vertices form a path (light red in Fig. 7); let them be numbered along this path, which we call the backbone of $G^{\prime}$. We assign every vertex $v$ that does not lie on the backbone to a neighboring backbone vertex; if $v$ is incident to an inner edge, we assign $v$ to the other endpoint of that edge. Otherwise $v$ has degree 2 and is incident to a unique backbone vertex via an outer edge, and we assign $v$ to this backbone vertex. For $i\in[k]$, backbone vertex $c_{i}$ induces, together with the vertices assigned to it, a fan $F_{i}$.
We draw the backbone in an x-monotone fashion. We start by drawing $F_{1}$ with the algorithm for drawing a fan as detailed in the proof of Lemma 1; see the leftmost gray box in Fig. 7. Then, for $i\in\{2,\dots,k\}$, we set $x(c_{i})$ to $x(c_{i-1})$ plus the number of inner edges incident to $c_{i}$ and we set $y(c_{i})$ depending (i) on the y-coordinates of the two neighbors of $c_{i}$ that have already been drawn ($c_{i-1}$ and the common neighbor $u_{i-1}$ of $c_{i-1}$ and $c_{i}$ in $F_{i-1}$) and (ii) on the directions of the edges that connect these vertices to $c_{i}$; see, for example, the placement of $c_{5}$ in Fig. 7. Then we draw $F_{i}$ with respect to the position of $c_{i}$, again using the algorithm from the proof of Lemma 1 with the following modifications. In general, vertices in $F_{i}$ that are adjacent to $c_{i}$ via an edge directed towards $c_{i}$ (resp. from $c_{i}$) are placed below (resp. above) all vertices in the drawings of $F_{1},\dots,F_{i}$; see the dark gray boxes below (resp. above) $c_{2},\dots,c_{5}$ in Fig. 7. If an edge of $F_{i}$ connects two neighbors of $c_{i}$ one of which lies above $c_{i}$ and one of which lies below $c_{i}$, then we route this edge to the left of all drawings of $F_{1},\dots,F_{i-1}$.
An exception to this rule occurs if $c_{i}$ and the common neighbor $w_{i}$ of $c_{i-1}$ and $c_{i}$ in $F_{i}$ must be both above or both below $c_{i-1}$ due to the directions of the corresponding edges. Let $u_{i-1}$ be the common neighbor of $c_{i-1}$ and $c_{i}$ in $F_{i-1}$. We assume, without loss of generality, that $c_{i}$ is above $c_{i-1}$. Let $P_{1}$ and $P_{2}$ be the first and second maximal subpath from Lemma 1 applied to $F_{i}$, and let $e_{i}$ be the edge connecting $P_{1}$ and $P_{2}$. We distinguish two subcases.
If $P_{1}$ is a directed path leaving $w_{i}$, then we draw $P_{1}$ above the edge $c_{i-1}c_{i}$ and we draw the edge $e_{i}$ straight, without going around all drawings of $F_{1},\dots,F_{i-1}$. In this case $e_{i}$ is directed from $P_{1}$ to $P_{2}$. Hence, $e_{i}$ crosses the edge $u_{i-1}c_{i}$ if $u_{i-1}c_{i}$ is directed from $c_{i}$ to $u_{i-1}$; see the situation for $c_{2}$ in Fig. 7. Note that by Property 4 of Lemma 1, $e_{i}$ may receive at most a second crossing when we draw the remainder of $F_{i}$ in the usual way.
Otherwise, that is, if $P_{1}$ contains an edge directed towards the first endpoint $w_{i}$ of $P_{1}$, let $f$ be the first such edge. We then place the part of $P_{1}$ up to the first endpoint of $f$ below the edge $c_{i-1}c_{i}$; see $w_{5}$ and $f$ in Fig. 7. If the edge $u_{i-1}c_{i}$ is directed towards $c_{i}$, we draw it between $w_{i}$ and the edge $c_{i-1}c_{i}$. Then it crosses the edge $c_{i-1}w_{i}$ but no other edge. We place the second endpoint of $f$ below all vertices in $V(F_{1})\cup\dots\cup V(F_{i-1})$ and continue with the remainder of $F_{i}$ as usual.
In any case, if $1<i<k$, then the last vertex $u_{i-1}$ of $F_{i-1}$ is connected to $c_{i}$ and $c_{i-1}$ is connected to the first vertex $w_{i}$ in $F_{i}$. These two edges may cross each other; see the crossings highlighted in orange in Fig. 7. If the edge $c_{i-1}w_{i}$ goes, say, up but the following outer edges go down until a vertex $v_{k}$ below $c_{i}$ is reached, then the edge $c_{i-1}w_{i}$ may be crossed a second time by the edge $v_{k-1}v_{k}$; see the crossing labeled $x$ on the edge $c_{3}w_{4}$ in Fig. 7. Observe that in this case the path $P_{1}$ is the directed path from $w_{i}$ to $v_{k-1}$. Thus, due to Property 4 of Lemma 1, edge $v_{k-1}v_{k}$ had been crossed at most once within its fan. Also $c_{i-1}w_{i}$ cannot have a third crossing. Thus, all edges are crossed at most twice. ∎
One can argue that every maximal pathwidth-2 graph can be generated from a maximal outerpath by connecting some pairs of adjacent vertices using an arbitrary number of (new) paths of length 2. In spite of the simplicity of this operation, we cannot hope to generalize the above result to pathwidth-2 graphs; see the linear lower bound on the upward local crossing number for such graphs stated in Theorem 2.
5 Testing Upward 1-Planarity
[…]
We start with a definition. Let $G_{1}$ and $G_{2}$ be any two st–digraphs. Let $s_{i}$ be the source and $t_{i}$ be the sink of $G_{i}$, with $i=1,2$. Let $G$ be a digraph that contains both $G_{1}$ and $G_{2}$ as induced subgraphs. Let $\Gamma$ be a drawing of $G$ and let $\Gamma_{1,2}$ be the drawing obtained by restricting $\Gamma$ to the vertices and edges of $G_{1}\cup G_{2}$. We say that $G_{1}$ and $G_{2}$ fully cross in $\Gamma$ if in $\Gamma_{1,2}$ every $s_{1}t_{1}$-path (i.e., a path directed from $s_{1}$ to $t_{1}$) crosses every $s_{2}t_{2}$-path. See Figs. 8(a) and 8(b) for examples of st-digraph $G_{1}$ and $G_{2}$ that do not fully cross or fully cross in a drawing of $\Gamma_{1,2}$, respectively.
Figure 8: Illustrations for the definition of fully crossing st-subgraphs. (a) and (b) Two st-digraphs $G_{1}$ and $G_{2}$ that do not fully cross, as witnessed by the two non-crossing dashed paths. (c) Two st-digraphs $G_{1}$ and $G_{2}$ that fully cross.
We now define a few gadgets; all of them are planar st-graphs. For positive integers $b$ and $q$, let a $(b,q)$-parallel be the parallel composition of $b$ oriented paths each consisting of $q$ edges; see Fig. 9(a). For a positive integer $p$, let a $(p)$-gate be the parallel composition of an oriented edge and a $(p-1,2)$-parallel; see Fig. 9(b). For positive integers $h$, $q$, and $a$, let an $(h,q,a)$-chain consist of a series of $h$ $(q)$-gates, followed by exactly one $(a)$-gate, followed again by $h$ $(q)$-gates; see Fig. 9(c).
(a) a $(4,3)$-parallel
(b) a $(4)$-gate
(c) a $(2,5,3)$-chain
Figure 9: Illustrations for the gadgets used in the construction of $G_{A}$ and of $G_{B}$.
[…]
We associate with a given instance $I$ of 3-Partition two planar st-graphs $G_{A}$ and $G_{B}$ defined as follows. Digraph $G_{A}$ is the parallel composition of $(b-1,W+1,a_{i})$-chains, one for every $i\in\{1,\dots,k\}$. Digraph $G_{B}$ is a $(b,q)$-parallel, with $q=W+(k-3)(W+1)$. Note that the underlying undirected graphs of both $G_{A}$ and $G_{B}$ are series-parallel.
Theorem 5.
Let $I$ be an instance of 3-Partition and let $G_{A}$ and $G_{B}$ be the two planar st-graphs associated with $I$. Assume there exists a digraph $G$ containing $G_{A}$ and $G_{B}$ as subgraphs with the following two properties: (i) if $G$ is upward 1-planar then $G_{A}$ fully crosses $G_{B}$ in every upward 1-planar drawing of $G$; (ii) if there exists an upward 1-planar drawing of the union of $G_{A}$ and $G_{B}$ in which $G_{A}$ and $G_{B}$ fully cross, then there exists an upward 1-planar drawing of $G$. Then the digraph $G$ is upward 1-planar if and only if $I$ admits a solution.
Proof.
Assume that $G$ is upward 1-planar and let $\Gamma$ be any upward 1-planar drawing of $G$. We prove that $\Gamma$ provides a solution of instance $I$ of 3-Partition. By hypothesis (i), $G_{A}$ and $G_{B}$ fully cross in $\Gamma$. Observe that only one path among the $b$ paths of the $(b,q)$-parallel $G_{B}$ can traverse the $(a_{i})$-gate of a $(W+1,b-1,a_{i})$-chain of $G_{A}$, where $a_{i}$, with $i=1,\dots,k$, is an element of the instance $I=\{a_{1},a_{2},a_{3},\dots a_{k}\}$ of 3-Partition, as otherwise the directed edge connecting the source and the sink of the $(a_{i})$-gate would be traversed more than once. Since $G_{A}$ and $G_{B}$ fully cross in $\Gamma$, by definition every path of $G_{B}$ crosses all the $(W+1,b-1,a_{i})$-chains of $G_{A}$. In particular, every path of $G_{B}$ must cross at least three $(a_{i})$-gates. Indeed, if a path $\pi$ of $G_{B}$ crossed less than three $(a_{i})$-gates of $G_{A}$, it would cross at least $k-2$ of the $(W+1)$-gates in the $(W+1,b-1,a_{i})$-chains of $G_{A}$. In order to have at most one crossing per edge, path $\pi$ should have at least $(k-2)(W+1)$ edges; however, by construction, $\pi$ has $W+(k-3)(W+1)=(k-2)(W+1)-1$ edges. Also, observe that by definition of $G_{A}$, if one path of $G_{B}$ crossed more than three $(a_{i})$-gates, some other path of $G_{B}$ should cross at most two $(a_{i})$-gates. Therefore, every path $\pi$ of $G_{B}$ must cross exactly three $(a_{i})$-gates and $k-3$ of the $(W+1)$-gates in $\Gamma$. Since every path $\pi$ of $G_{B}$ uses $(k-3)(W+1)$ of its edges to cross $k-3$ of the $(W+1)$-gates in $\Gamma$, $\pi$ can cross at most $W$ edges of the $(a_{i})$-gates. Since $\sum_{i=1}^{k}a_{i}=W\cdot b$, the number of crossings of each $\pi$ with its three $(a_{i})$-gates must be exactly $W$. It follows that if $G$ has an upward 1-planar drawing then the instance $I$ of 3-Partition admits a solution.
Conversely, assume that the instance $I$ of 3-Partition admits a solution. In order to prove that $G$ admits an upward 1-planar drawing, by hypothesis (ii) it suffices to construct an upward 1-planar drawing $\Gamma_{AB}$ of $G_{A}$ and $G_{B}$ where $G_{A}$ and $G_{B}$ fully cross. We enumerate the paths of $G_{B}$ as $\pi_{1},\pi_{2},\dots,\pi_{b}$. We also enumerate the $2(b-1)$ many $(W+1)$-gates and the $(a_{i})$-gate of a $(b-1,W+1,a_{i})$-chain as $g_{i,1},g_{i,2},\cdots,g_{i,2(b-1)}$, where $g_{i,1}$ contains the source $s_{A}$ of $G_{A}$ and $g_{i,2(b-1)}$ contains the sink $t_{A}$ of $G_{A}$ (see also Fig. 10). Consider a path $\pi_{j}$ with $1\leq j\leq b$, and let $\{a_{\chi},a_{\lambda},a_{\mu}\}$ be the $j$-th bin of the solution of $I$. Let $\nu$ be an index such that $1\leq\nu\leq b$.
Figure 10: (a) Digraph $G_{A}$ (dark red) and a schematic representation of digraph $G_{B}$ where each colored curve represents a directed path with $W+(k-3)(W+1)$ edges. The corresponding instance of 3-Partition is $I=\{1,1,1,2,2,2,2,3,4\}$, with $b=3$ and $W=6$. The 1-planar drawing corresponds to the solution $\{1,1,4\}$ (green path), $\{2,2,2\}$ (blue path), and $\{1,2,3\}$ (red path). The drawing in (a) is not upward but it can be made upward by stretching it vertically as shown in (b), where thick edges represent $(q)$-gates and the central white-filled edges represent $(a)$-gates.
  • If $\nu$ is one of $\chi,\lambda,\mu$, then $\pi_{j}$ crosses the $(a_{\nu})$-gate of the $(b-1,W+1,a_{\nu})$-chain (see, for example, the red path crossing $g_{1,3}$ in Fig. 10).
  • If $\nu$ is not one of $\chi,\lambda,\mu$ and the path $\pi_{h}$ that crosses the $(a_{\nu})$-gate of the $(b-1,W+1,a_{\nu})$-chain is such that $h>j$, then $\pi_{j}$ crosses the gate $g_{\nu,j}$ of the $(W+1,b-1,a_{\nu})$-chain (see, for example, the blue path crossing $g_{2,2}$ in Fig. 10).
  • Otherwise ($h<j$), path $\pi_{j}$ crosses the gate $g_{\nu,b-1+j}$ of the $(b-1,W+1,a_{\nu})$-chain (see, for example, the red path crossing $g_{5,5}$ in Fig. 10).
By the procedure above, if $\pi_{j}$ crosses a gate $g(\nu,q)$, with $1\leq\nu\leq k$ and $1\leq q\leq 2(b-1)$, there is no other $\pi_{h}$, with $h\neq j$, such that $\pi_{h}$ crosses $g(\nu,q)$. Also, the number of edges that $\pi_{j}$ crosses is $a_{\chi}+a_{\lambda}+a_{\mu}+(k-3)(W+1)=W+(k-3)(W+1)$, which is the number of edges of $\pi_{j}$. Hence, it is possible to draw $\pi_{j}$ in such a way that each of its edges is crossed exactly once and no edge of $G_{A}$ is crossed more than once. This concludes the proof. ∎
Theorem 6.
[…]
  1. 1.
    The bimodal rotation system is fixed, the DAG has exactly one source and exactly one sink, the underlying graph is series-parallel.
  2. 2.
    The bimodal rotation system is not fixed, the DAG has exactly one source and exactly one sink, the underlying planar graph is obtained by replacing the edges of a $K_{4}$ with series-parallel graphs.
  3. 3.
    The bimodal rotation system is not fixed, the underlying graph is series-parallel, there is one source and two sinks.
Proof.
It is immediate to observe that upward 1-planarity testing is in the NP class of complexity (one can guess an upward embedding and test it in polynomial time). We show that it is NP-hard. For each case in the statement it suffices to exhibit a digraph $G$ that contains $G_{A}$ and $G_{B}$ as sub-graphs and that satisfies the conditions of Theorem 5.
Let $m_{A}$ and $m_{B}$ be the number of edges of $G_{A}$ and $G_{B}$, respectively. We define a barrier to be an st-digraph consisting of a $(d,2)$-parallel, where $d=m_{A}+m_{B}+1$. Note that no $s_{A}t_{A}$-path ($s_{B}t_{B}$-path) can fully cross a barrier in such a way that every edge of the path is crossed at most once. This implies that neither $G_{A}$ nor $G_{B}$ can fully cross a barrier.
Figure 11: Some digraphs for the proof of Theorem 6. Thick black edges represent barriers.
  • Case 1: Refer to Fig. 11(a), where the thick edges schematically represent barriers $B_{1}$ (from $s_{A}$ to $s_{B}$), $B_{2}$ (from $s_{B}$ to $t_{A}$), and $B_{3}$ (from $t_{A}$ to $t_{B}$); in the figure, the red shaded shape represents $G_{A}$ and the blue shaded shape represents $G_{B}$. The figure schematically represents a series-parallel digraph $G$ with source $s_{A}$ and sink $t_{B}$. As described in Fig. 11(b), an upward 1-planar drawing where the edges of $G_{A}$ precede the edges of $B_{1}$ in the left to right order of the edges exiting $s_{A}$ and the edges of $G_{B}$ precede the edges of $B_{3}$ in the left to right order of the edges entering $t_{B}$ exists if and only if $G_{A}$ crosses $G_{B}$. In fact, in any upward 1-planar drawing of $G$ that preserves its bimodal rotation system $G_{A}$ crosses $G_{B}$; conversely, starting from an upward 1-planar drawing of the union of $G_{A}$ and $G_{B}$ in which $G_{A}$ and $G_{B}$ cross, an upward 1-planar drawing of $G$ that preserves the bimodal rotation system of Fig. 11(a) can be obtained as shown in Fig. 11(b).
  • Case 2: We modify a very well known small digraph that is traditionally used to exhibit a digraph that is planar but not upward planar (Bertolazzi, Di Battista, Mannino, and Tamassia, 1998; Di Battista, Eades, Tamassia, and Tollis, 1999). As above, in Fig. 11(c) thick edges schematically represent barriers, the red shaded shape represents $G_{A}$, and the blue shaded shape represents $G_{B}$. Again, an upward 1-planar drawing of this digraph exists if and only if $G_{A}$ fully crosses $G_{B}$ (see Fig. 11(d)). In particular, neither $s_{B}$ nor $t_{B}$ can be drawn in the finite region of the plane bounded by $G_{A}$ since at least one $s_{A}t_{A}$-path would fully cross a barrier. Similarly, neither $s_{A}$ nor $t_{A}$ can be drawn in the finite region of the plane bounded by $G_{B}$. Observe that the digraph has a single source and a single sink and its underlying planar graph is obtained by replacing the edges of a $K_{4}$ with series-parallel graphs.
  • Case 3: Consider the digraph $G$ of Fig. 11(e), where the thick edges schematically represent barriers; as in the previous case, the red shaded shape represents $G_{A}$ and the blue shaded shape represents $G_{B}$. With analogous arguments as in the previous case, it is immediate to see that $G$ has an upward 1-planar drawing if and only if $G_{A}$ fully crosses $G_{B}$ (see Fig. 11(f)).
The following corollary is a consequence of the argument used to prove the second case in the statement of Theorem 6.
Corollary 3.
[…]
Table 1: A comparison between results in the literature about the complexity of testing upward planarity and the results discussed in this paper about the complexity of testing upward 1-planarity.
Underlying Acyclic Upward planarity Upward 1-planarity
planar orientation: Embedding Rotation system
graph sources/sinks fixed variable fixed variable
[…]
single/single polynomial (Chaplick et al., 2022a, b; Didimo et al., 2010) trivially polynomial
[…]
single/single polynomial (Chaplick et al., 2022a)
6 Testing Upward Outer-1-Planarity
Theorem 6 motivates the study of the complexity of testing upward 1-planarity with additional constraints. A common restriction in the study of beyond-planar graph drawing problems is the one that requires that all vertices are incident to the same face. Specifically, a graph is upward outer-1-planar if it admits an upward 1-planar embedding in which every vertex lies on the outer face. This section is devoted to the proof of the following result.
Theorem 7.
For single-source DAGs, upward outer-1-planarity can be tested in linear time.
6.1 Basic Facts and Definitions
We provide the following characterization for the single-source DAGs admitting an upward outer-1-planar drawing.
Theorem 8.
A single-source graph is upward outer-1-planar if and only if it admits an outer-1-planar embedding whose planarization is acyclic.
Proof.
Let $G$ be a single-source graph with an upward outer-1-planar embedding $\Gamma$. Clearly, $\Gamma$ is outer-1-planar. Moreover, planarizing $\Gamma$ yields an upward drawing of the planarization, hence the planarization is acyclic.
Conversely, let $\Gamma$ be an outer-1-planar embedding whose planarization $\Gamma^{\star}$ is acyclic. Observe that no crossing vertex of $\Gamma^{\star}$ is a source or a sink. Therefore, $\Gamma^{\star}$ only has a single source $s$ and the sinks of $G$, which are all incident to the outer face. Let $\Gamma^{+}$ be obtained from $\Gamma^{\star}$ by adding a new super sink $t$, with edges from all sinks of $G$ to $t$ into the outer face. Clearly $\Gamma^{+}$ is acyclic and a planar $st$-graph, and therefore upward planar (Di Battista and Tamassia, 1988). Removing $t$ then yields an upward planar drawing of $\Gamma^{\star}$ and hence $\Gamma$ is upward outer-1-planar. ∎
The following lemma follows from Theorem 8 and shows that we can decompose the problem on the biconnected components of the input graph.
Lemma 2.
A single-source graph is upward outer-1-planar if and only if each of its blocks admits an upward outer-1-planar embedding.
Proof.
The necessity is trivial. For the sufficiency, suppose that all blocks admit an upward outer-1-planar embedding, whose planarization is thus acyclic. Combine all such planarizations at the cut-vertices so that all their vertices are incident to the outerface. Since each cycle in the graph is contained inside a block, this yields an acyclic planarization of an outer-1-planar embedding of the entire graph. This and Theorem 8 imply the statement. ∎
The SPQR-tree data structure, introduced by Di Battista and Tamassia (1996), to represent the decomposition of a biconnected graph into its triconnected components is a special type of decomposition tree (Hopcroft and Tarjan, 1973). A decomposition tree of a biconnected graph $G$ is a tree $\mathcal{T}$ whose nodes $\mu$ are associated with a biconnected multi-graph $\operatorname{skel}(\mu)$, called skeleton. The edges of a skeleton can be either real or virtual, and for each node $\mu$, the virtual edges of $\operatorname{skel}(\mu)$ correspond bijectively to the edges of $\mathcal{T}$ incident to $\mu$. The tree $\mathcal{T}$ can be inductively defined as follows. In the base case, $\mathcal{T}$ consists of a single node $\mu$ whose skeleton is the graph $G$ consisting solely of real edges. In the inductive case, let $\mu$ be any node of $\mathcal{T}$ and let $H_{1}$ and $H_{2}$ be edge-disjoint connected subgraphs of $\operatorname{skel}(\mu)$ such that $\operatorname{skel}(\mu)=H_{1}\cup H_{2}$ and $H_{1}\cap H_{2}=\{u,v\}$. We obtain a new decomposition tree $\mathcal{T}^{\prime}$ from $\mathcal{T}$ by splitting the node $\mu$ into two adjacent nodes $\nu_{1}$ and $\nu_{2}$ whose skeletons are $H_{1}+uv$ and $H_{2}+uv$, respectively, where $uv$ is a new virtual edge that corresponds to the edge $\nu_{1}\nu_{2}$ of $\mathcal{T}^{\prime}$; also, we replace each edge $\mu\tau$ of $\mathcal{T}$ either with the edge $\nu_{1}\tau$ or with the edge $\nu_{2}\tau$ depending on wether edge $\mu\tau$ corresponds to a virtual edge of $H$ that belongs to $H_{1}$ or to $H_{2}$, respectively. The edges of $\mathcal{T}$ incident to $\mu$ are distributed among $\nu_{1}$ and $\nu_{2}$ based on whether their corresponding virtual edge belongs to $H_{1}$ or $H_{2}$. Let now $\mathcal{T}$ be an arbitrary decomposition tree of a biconnected graph. Note that each edge of $G$ occurs as a real edge in precisely one skeleton of $\mathcal{T}$. Consider a virtual edge $e$ in the skeleton of some node $\mu$ of $\mathcal{T}$ and let $\varepsilon$ be the edge of $\mathcal{T}$ which corresponds to $e$. The expansion graph $\operatorname{expn}(e)$ of $e$ is the graph obtained as the union of the real edges belonging to the skeletons of the subtree of $\mathcal{T}$ reachable from $\mu$ via the edge $\varepsilon$. Observe that $\operatorname{expn}(\{a,b\})$ is connected to the rest of the graph through $a$ and $b$. The vertices $a$ and $b$ are the poles of $\operatorname{expn}(\{a,b\})$. For every pair of adjacent nodes $\mu$ and $\nu$ of $\mathcal{T}$, there exists a virtual edge $e$ in $\operatorname{skel}(\mu)$ and a virtual edge $e^{\prime}$ in $\operatorname{skel}(\nu)$ with the same end-vertices. We say that $\mu$ (resp. $\nu$) is the refining node $\operatorname{refn}(e^{\prime})$ of $e^{\prime}$ (resp. $\operatorname{refn}(e)$ of $e$).
The SPQR-tree $\mathcal{T}$ of a biconnected graph $G$ is a decomposition tree that has three types of nodes. The skeleton of an S-node is a simple cycle of length at least 3, the skeleton of a P-node consists of two vertices and at least three parallel edges, and the skeleton of an R-node is a simple 3-connected graph. Moreover, no two S-nodes and no two P-nodes are adjacent in $\mathcal{T}$. We remark that unlike the classical definition of SPQR-trees, we allow real edges in any skeleton; this avoids the need to use Q-nodes whose skeletons have two vertices and one real and one virtual edge.
The SPQR-tree of a biconnected planar graph $G$ can be used to succinctly represent all planar embeddings of $G$. Specifically, any planar embedding of $G$ uniquely defines an embedding for all the skeletons of $\mathcal{T}$. Moreover, recursively combining planar embeddings of the skeletons of $\mathcal{T}$ via 2-clique sums of such embeddings results in a planar embedding of $G$. The 2-clique-sum of two embeddings $\mathcal{E}^{\prime}$ and $\mathcal{E}^{\prime\prime}$ containing the virtual edge $\{u,v\}$ identifies the two copies of $u$ and the two copies of $v$, removing the edge $\{u,v\}$, in such a way that edges in $\mathcal{E}^{\prime}$ do not alternate with edges in $\mathcal{E}^{\prime\prime}$. Even in the presence of non-planar embeddings of skeletons, such an operation can be performed in the same fashion as long as the common virtual edge $\{u,v\}$ is crossing free in both embeddings.
Let $\mu$ be a P-node with skeleton vertices $\{u,v\}$, let $e$ be a virtual edge in $\operatorname{skel}(\mu)$ that is refined by an S-node $\lambda$ with skeleton $(u,c_{1},c_{2},\dots,c_{k},v,u)$. Then the first segment of $e$ at $u$ is the (virtual or real) edge $\{u,c_{1}\}$ of $\operatorname{skel}(\lambda)$ and the first segment of $e$ at $v$ is the (virtual or real) edge $\{c_{k},v\}$ of $\operatorname{skel}(\lambda)$.
6.2 Proof of Theorem 7
There exist two different algorithms for testing outer-1-planarity, namely by Hong, Eades, Katoh, Liotta, Schweitzer, and Suzuki (2013, 2015) and by Auer, Bachmaier, Brandenburg, Gleißner, Hanauer, Neuwirth, and Reislhuber (2016). It is likely that both can be adapted to test upward outer-1-planarity. We choose to follow the approach by Auer et al. (2016), which we briefly summarize in the following, as it seems slightly simpler to extend to our setting. Two key properties that both papers leverage are the facts that if $G$ is an outer-1-planar graph and $\mathcal{T}$ is its SPQR-tree, then the skeleton of each R-node of $\mathcal{T}$ is a $K_{4}$ that contains two non-adjacent real edges and the skeleton of each P-node has at most four virtual edges and at most one real edge. In what follows, we assume that these two properties are satisfied.
6.2.1 The Outer-1-Planarity Testing Algorithm of Auer et al.
By Lemma 2, we may assume that our input graph $G$ is biconnected and acyclic. Consider a decomposition tree $\mathcal{T}$ of $G$ (not necessarily the SPQR-tree). Let $\mu$ be a node of $\mathcal{T}$. We call an outer-1-planar embedding of $\operatorname{skel}(\mu)$ good if only real edges cross and each virtual edge is incident to the outer face.
Observation 1.
Good embeddings of all skeletons of $\mathcal{T}$ together define a unique outer-1-planar embedding of $G$, which is obtained by forming the 2-clique sums of the embeddings of the skeletons.
However, if we consider the SPQR-tree, good embeddings of the skeletons may not exist even if $G$ is outer-1-planar. Therefore, Auer et al. (2016) work with a more general definition of outer-1-planar embeddings for skeletons of the SPQR-tree and show that an outer-1-planar embedding of $G$ can be combinatorially described by embeddings of the skeletons of its SPQR-tree $\mathcal{T}$ that satisfy the following five conditions C1–C5. We note that the edges of the skeletons may cross even for P-nodes, i.e., the embeddings of the skeletons we consider are not necessarily simple drawings.
  1. C1
    Each skeleton is embedded outer-1-planar such that each virtual edge has a segment that is incident to the outer face (Auer et al., 2016, Proposition 3).
  2. C2
    Skeletons of S-nodes are embedded planarly (Auer et al., 2016, Corollary 3).
  3. C3
    Skeletons of R-nodes are embedded such that virtual edges are crossing-free and precisely two non-virtual edges cross each other (Auer et al., 2016, Corollary 2).
  4. C4
    If an edge $e=\{u,v\}$ of the skeleton of a P-node is crossed, then it is a virtual edge corresponding to an S-node. Moreover, if the segment of $e$ incident to $u$ (incident to $v$) is not incident to the outer face, then the first segment of $e$ at $u$ (at $v$) must be a real edge (Auer et al., 2016, Lemma 2).
  5. C5
    Let $\mu$ be an S-node whose skeleton contains a real edge $\{u,v\}$ and two adjacent virtual edges $\{u^{\prime},u\}$ and $\{v,v^{\prime}\}$ that are refined by $\rho$ and $\lambda$, respectively. Let $e_{\rho},e_{\lambda}$ be the virtual edges that represent $\mu$ in $\operatorname{skel}(\rho)$ and $\operatorname{skel}(\lambda)$, respectively. If both $\rho$ and $\lambda$ are P-nodes, then the segment of $e_{\rho}$ incident to $u$ must be incident to the outer face in the embedding of $\operatorname{skel}(\rho)$ or the segment of $e_{\lambda}$ incident to $v$ must be incident to the outer face in the embedding of $\operatorname{skel}(\lambda)$.
Figure 12: A portion of an SPQR-tree with an S-node $\mu$ and two adjacent S-nodes $\rho,\lambda$ illustrating the necessity of Condition 5. Since neither the segment of $e_{\rho}$ incident to $u$ is incident to the outer face in the embedding of $\operatorname{skel}(\rho)$ nor the segment of $e_{\lambda}$ incident to $v$ is incident to the outer face in the embedding of $\operatorname{skel}(\lambda)$, the real edge $\{u,v\}$ must be involved in at least two crossings. Namely, at least one crossing with an edge in the expansion graph of the blue virtual edges of $\operatorname{skel}(\rho)$ and at least one crossing with an edge in the expansion graph of the yellow virtual edges of $\operatorname{skel}(\lambda)$.
The idea behind Conditions 4 and 5 is that crossing two virtual edges $e,e^{\prime}$ in a skeleton of a P-node $\mu$ with vertices $\{u,v\}$ in such a way that $e$ is not incident to the outer face at $u$ and $e^{\prime}$ is not incident to the outer face at $v$ implies that we cross the first segment of $e$ at $u$ and the first segment of $e^{\prime}$ at $v$; see the gray-shaded regions in Fig. 12 and the more formal discussion later. Condition 4 guarantees that these are real edges. On the other hand, Condition 5 ensures that each real edge receives at most one crossing in this way; see Fig. 12 for an illustration of a case in which this is violated.
It follows from the work of Auer et al. (2016) that together these conditions describe all outer-1-planar embeddings of $G$ without unnecessary crossings (i.e., there is no outer-1-planar drawing whose crossing edge pairs form a strict subset). We summarize this in the following theorem. As this result does not appear explicitly in the work of Auer et al., for the sake of completeness, we briefly sketch the construction. In particular, the construction of an embedding of $G$ from the embeddings of the skeletons is crucial for understanding our next steps.
Theorem 9.
Let $G$ be a biconnected graph and let $\mathcal{T}$ be its SPQR-tree. There is a bijection between the outer-1-planar embeddings of $G$ without unnecessary crossings and the choice of an embedding for each skeleton of $\mathcal{T}$ that satisfy Conditions 15. Moreover, both directions of the bijection can be computed in linear time.
Proof.
For the bijection, observe that the necessity of 15 has been argued above. For the converse, we show how we obtain an embedding of $G$ from embeddings $\{\mathcal{E}_{\mu}\}_{\mu\in V(\mathcal{T})}$ for each skeleton $\operatorname{skel}(\mu)$ that satisfy 15. We refer to this embedding as the combined embedding of $\{\mathcal{E}_{\mu}\}_{\mu\in V(\mathcal{T})}$.
Figure 13: Illustration for the extension of a P-node $\mu$.
First note that, if there are no crossings that involve virtual edges in the embeddings of the skeletons, then the embeddings are good and their 2-clique sum yields an outer-1-planar embedding of $G$ by 1. If there are virtual edges that are involved in a crossing, they belong to the skeletons of P-nodes by Conditions 2 and 3. We reduce to the case of good embeddings by extending such P-nodes, that is, by modifying the SPQR-tree, its skeletons, and their embeddings as follows; refer to Fig. 13. Let $e^{\prime},e^{\prime\prime}$ be two crossing virtual edges in an embedding of a skeleton of a P-node $\mu$. Let $\{u,v\}$ be the vertices of $\operatorname{skel}(\mu)$. By Condition 4, we have that the nodes $\nu^{\prime}=\operatorname{refn}(e^{\prime})$ and $\nu^{\prime\prime}=\operatorname{refn}(e^{\prime\prime})$ are S-nodes with skeletons $(u,c^{\prime}_{1},\dots,c^{\prime}_{k},v,u)$ and $(u,c_{1}^{\prime\prime},\dots,c_{\ell}^{\prime\prime},v,u)$, respectively. Without loss of generality, we may assume that the segment of $e^{\prime}$ incident to $u$ is incident to the outer face in the embedding of $\operatorname{skel}(\mu)$, and hence by Condition 1 the segment of $e^{\prime\prime}$ incident to $v$ is also incident to the outer face. By Condition 4, the edge $\{c_{k}^{\prime},v\}$ of $\operatorname{skel}(\nu^{\prime})$, which is the first segment of $e^{\prime}$ at $v$, and the edge $\{u,c_{1}^{\prime\prime}\}$ of $\operatorname{skel}(\nu^{\prime\prime})$, which is the first segment of $e^{\prime\prime}$ at $u$, are real edges. We perform the following modifications. First, we shorten the skeletons of $\nu^{\prime}$ and of $\nu^{\prime\prime}$ to $(u,c^{\prime}_{1},\dots,c^{\prime}_{k-1},c^{\prime}_{k},u)$ and $(c^{\prime\prime}_{1},c^{\prime\prime}_{2},\dots,c^{\prime\prime}_{\ell},v,c^{\prime\prime}_{1})$, respectively, and keep their embeddings planar. Second, we subdivide the virtual edge $e^{\prime}$ ($e^{\prime\prime}$) with the vertex $c^{\prime}_{k}$ ($c^{\prime\prime}_{1}$) such that $\{v,c^{\prime}_{k}\}$ ($\{u,c^{\prime\prime}_{1}\}$) is a real edge and $\{c^{\prime}_{k},u\}$ ($\{c^{\prime\prime}_{1},v\}$) is a virtual edge that is refined by the modified S-node $\nu^{\prime}$ ($\nu^{\prime\prime}$). We embed the skeletons of such nodes so that the real edges $\{v,c^{\prime}_{k}\}$ and $\{u,c^{\prime\prime}_{1}\}$ cross and the two virtual edges $\{u,c^{\prime}_{k}\}$ and $\{v,c^{\prime\prime}_{1}\}$ are incident to the outer face. Note that this essentially moves two real edges from S-node skeletons into an adjacent P-node. See the right side of Fig. 13 for the modified skeletons.
Condition 5 guarantees that no real edge needs to be moved into the skeletons of two different P-nodes. Hence we can apply this operation simultaneously and independently to all pairs of crossing virtual edges. We then arrive at a decomposition tree $\mathcal{T}^{*}$ of $G$, whose skeletons have good embeddings. As explained above, we then obtain an outer-1-planar embedding of $G$.
Observe that this construction can be carried out in linear time. Conversely, given an outer-1-planar embedding $\mathcal{E}$ of $G$. The embeddings $\mathcal{E}_{\mu}$ of all skeletons $\operatorname{skel}(\mu)$ can be obtained by checking the order of the incident edges for each vertex and the pairs of edges that cross. Moreover, it is then clear that applying the above construction to the $\mathcal{E}_{\mu}$, we reobtain $\mathcal{E}$. ∎
The tree $\mathcal{T}^{\star}$ we constructed towards the end of the proof of Theorem 9 is called the extension of $\mathcal{T}$ with respect to $\{\mathcal{E}_{\mu}\}_{\mu\in V(\mathcal{T})}$. Since the modification neither adds nor removes nodes of $\mathcal{T}$, there is a bijection between the nodes of $\mathcal{T}$ and $\mathcal{T}^{*}$. Namely, each node $\mu$ of $\mathcal{T}$ contributes a unique node $\mu^{*}$ to $\mathcal{T}^{*}$, and these are all the nodes of $\mathcal{T}^{*}$. The node $\mu^{*}$ is called the extended node of $\mu$ and the good embedding of $\operatorname{skel}(\mu^{*})$ obtained with the procedure described above is called the extended embedding of $\operatorname{skel}(\mu)$. We further note that for P- and R-nodes $\mu$$\operatorname{skel}(\mu^{\star})$ only depends on the embedding $\mathcal{E}_{\mu}$ of $\operatorname{skel}(\mu)$. Thus, for P-nodes and R-nodes $\operatorname{skel}(\mu^{\star})$ can also be defined if we only have an embedding of $\operatorname{skel}(\mu)$ that satisfies 14. Condition 5 is only needed if we want to combine such skeletons into the whole extension $\mathcal{T}^{\star}$.
6.2.2 An Algorithm to test Upward Outer-1-Planarity.
Our goal is to find an outer-1-planar embedding of $G$ whose planarization is acyclic. Observe that it suffices to consider outer-1-planar embeddings without unnecessary crossings, since adding more crossings cannot make a cyclic planarization acyclic. Our goal is therefore to find outer-1-planar embeddings of the skeletons of the SPQR-tree $\mathcal{T}$ of $G$ satisfying Conditions 15 so that the resulting planarization of $G$ is acyclic.
Directing virtual edges.
To keep track of the existence of directed paths, we orient some edges of the skeletons of $\mathcal{T}$. Consider a skeleton $\operatorname{skel}(\mu)$ and let $e=\{u,v\}$ be an edge of $\operatorname{skel}(\mu)$. We orient $e$ from $u$ to $v$ if either it is a real edge directed from $u$ to $v$ in $G$ or if $\operatorname{expn}(e)$ contains a directed path from $u$ to $v$.
We observe some properties of the constructed orientation. First, note that a virtual edge $\{u,v\}$ is oriented in at most one direction, but it may also remain undirected in case its expansion graph contains neither a directed path from $u$ to $v$ nor a directed path from $v$ to $u$. Second, if a virtual edge $e$ with endpoints $u,v$ is directed, say from $u$ to $v$, then $u$ is the single source of $\operatorname{expn}(e)$, as otherwise $G$ would contain an additional source. Finally, if $e$ is undirected, since $G$ has a single source $s$, either $\operatorname{expn}(e)$ contains $s$ as a non-pole vertex or both $u$ and $v$ are sources of $\operatorname{expn}(e)$.
Acyclic Embeddings.
Figure 14: Two embeddings of the skeleton of a P-node $\mu$ and the corresponding planarizations. The planarization in (a) contains a directed cycle, the one in (b) does not. Thick arrowed edges show the direction of the virtual edges; a double arrow indicates an undirected virtual edge.
Suppose that the edges of the skeletons are directed as described above, let $\mu$ be a node with an embedding $\mathcal{E}_{\mu}$ that satisfies C1–C4. Recall from the remark after the proof of Theorem 9 that $\operatorname{skel}(\mu^{\star})$ and its embedding $\mathcal{E}_{\mu}^{\star}$ depend only on $\mathcal{E}_{\mu}$. We call $\mathcal{E}_{\mu}$ acyclic if the planarization of the extended embedding $\mathcal{E}_{\mu}^{\star}$ is acyclic; see Fig. 14 for examples. It is not surprising that indeed such embeddings are necessary for the existence of an upward outer-1-planar embedding of $G$.
Lemma 3 (Necessity).
Let $\mathcal{E}$ be an upward outer-1-planar embedding of $G$ without unnecessary crossings. Then the embedding of $\operatorname{skel}(\mu)$ induced by $\mathcal{E}$ is acyclic for each P- or R-node $\mu$ of the SPQR-tree $\mathcal{T}$ of $G$.
Proof.
By Theorem 9, $\mathcal{E}$ induces embeddings $\mathcal{E}_{\mu}$ for each skeleton $\operatorname{skel}(\mu)$ of $\mathcal{T}$, which satisfy the Conditions 15. Let $\mathcal{T}^{*}$ be the extension of $\mathcal{T}$ with respect to these embeddings and recall from the proof of Theorem 9 that $\mathcal{E}$ can be obtained as the 2-clique sum of the embeddings $\mathcal{E}_{\mu}^{\star}$. Let $\mu$ be a P- or R-node of $\mathcal{T}$ with embedding $\mathcal{E}_{\mu}$ and let $\mu^{*}$ be the extended node in $\mathcal{T}^{*}$ with embedding $\mathcal{E}_{\mu}^{*}$. As the only crossings in $\mathcal{E}_{\mu}^{*}$ are between real edges, we can form its planarization by replacing each crossing by a dummy vertex, thereby subdividing the involved directed edges into paths of length 2. Since the upward embedding $\mathcal{E}$ in particular induces an upward embedding of its planarization, it follows that the planarization of $\mathcal{E}_{\mu}^{\star}$ does not contain a cycle that consists of directed edges, i.e., $\mathcal{E}_{\mu}$ is acyclic. ∎
Lemma 4 (Sufficiency).
Let $\{\mathcal{E}_{\mu}\}_{\mu\in V(\mathcal{T})}$ be embeddings of the skeletons of $\mathcal{T}$ that satisfy Conditions 15. If $\mathcal{E}_{\mu}$ is acyclic for all P- and R-nodes, then the combined embedding of $\{\mathcal{E}_{\mu}\}_{\mu\in V(\mathcal{T})}$ is an outer-1-planar embedding of $G$ whose planarization is acyclic.
Proof.
Let $\mathcal{T}^{\star}$ be the extension of $\mathcal{T}$ with respect to the embeddings $\{\mathcal{E}_{\mu}\}_{\mu\in V(\mathcal{T})}$ of the skeletons and let $\mathcal{E}$ denote the combined embedding of $\{\mathcal{E}_{\mu}\}_{\mu\in V(\mathcal{T})}$. By Theorem 9, $\mathcal{E}$ is an outer-1-planar embedding of $G$. In the remainder, we show that its planarization $G^{\prime}$ is acyclic. Thus, assume for the sake of contradiction that $G^{\prime}$ contains a directed cycle $C$.
We will consider the projection of $C$ to the planarizations of the skeletons of the nodes of $\mathcal{T}^{\star}$. Let $\mu^{\star}$ be a node of $\mathcal{T}^{\star}$ and consider its planarization $\operatorname{skel}^{+}(\mu^{\star})$. An edge $e$ of $\operatorname{skel}^{+}(\mu^{\star})$ belongs to the projection of $C$ if either (i) $e$ is a real edge that belongs to $C$ or (ii) $e$ is a virtual edge whose expansion graph contains a real edge that belongs to $C$. Observe that for each planarized skeleton, the projection of $C$ is either a cycle $C^{\prime}$ or a single edge. In the former case, we also say that $C$ projects to the cycle $C^{\prime}$.
Claim: There exists a P- or R-node $\mu^{\star}$ of $\mathcal{T}^{\star}$ where the cycle $C$ projects to a cycle $C^{\prime}$ in the planarization $\operatorname{skel}^{+}(\mu^{\star})$ such that the source $s$ of $G$ is either a vertex of $\operatorname{skel}^{+}(\mu^{\star})$ or belongs to the expansion graph of an edge of $\operatorname{skel}^{+}(\mu^{\star})$ that does not belong to $C^{\prime}$.
To prove the claim, observe that the skeletons where $C$ projects to a cycle form a subtree $\mathcal{T}_{C}$ of $\mathcal{T}^{\star}$. Similarly, the nodes whose skeletons contain $s$ as a vertex also form a subtree $\mathcal{T}_{s}$ of $\mathcal{T}^{\star}$. We choose $\mu^{\star}$ either as a common node of these two subtrees (if one exists) or as the node of $\mathcal{T}_{C}$ that is closest to the subtree $\mathcal{T}_{s}$. In the former case, we have that $s$ is a vertex of $\operatorname{skel}^{+}(\mu^{\star})$. Note that, since $s$ is a source, it is not contained in $C$ and therefore it is also not contained in the projection of $C$ to $\operatorname{skel}^{+}(\mu^{\star})$. In particular $\mu^{\star}$ cannot be an S-node. In the latter case let $e$ denote the virtual edge of $\operatorname{skel}^{+}(\mu^{\star})$ whose expansion graph contains $s$. If $e$ is contained in the projection of $C$ to $\operatorname{skel}^{+}(\mu^{\star})$, then $\operatorname{refn}(e)$ belongs to $\mathcal{T}_{C}$ and is closer to $\mathcal{T}_{s}$ than $\mu^{\star}$, which contradicts the choice of $\mu^{\star}$. Note that this in particular excludes the case that $\mu^{\star}$ is an S-node. This completes the proof of the claim.
Consider the node $\mu^{\star}$ from the claim and let $\mu$ be the corresponding node of $\mathcal{T}$. The projection $C^{\prime}$ of $C$ to $\operatorname{skel}^{+}(\mu^{\star})$ may contain both real and virtual edges. If all edges of $C^{\prime}$ are directed, the cycle is already present in $\operatorname{skel}^{+}(\mu^{\star})$, i.e., the embedding $\mathcal{E}_{\mu}$ of $\operatorname{skel}(\mu)$ is not acyclic; a contradiction. Since real edges are always directed, it hence follows that $C^{\prime}$ contains an undirected virtual edge $\{u,v\}$ of $\operatorname{skel}(\mu^{\star})$. Since in expanded skeletons only real edges have crossings, the edge $\{u,v\}$ is uncrossed in the embedding of $\operatorname{skel}(\mu^{\star})$. Therefore $\operatorname{expn}(\{u,v\})$ is not crossed by edges outside of $\operatorname{expn}(\{u,v\})$. In particular the planarization of $\operatorname{expn}(\{u,v\})$ is connected to the rest of the graph only via the vertices $u$ and $v$ in the planarization $G^{\prime}$ of $G$. Since $\{u,v\}$ belongs to $C^{\prime}$, the planarization of $\operatorname{expn}(\{u,v\})$ contains a directed path $\pi$ between its two poles, say from $u$ to $v$. However, since $\{u,v\}$ is undirected and does not contain $s$, both $u$ and $v$ are sources in $\operatorname{expn}(\{u,v\})$ and hence they are sources also in its planarization. This contradicts the existence of $\pi$, and hence proves the lemma. ∎
In light of Lemmas 3 and 4, it suffices to test whether the skeletons of $\mathcal{T}$ admit embeddings that satisfy 15 and that are acyclic for P- and R-nodes. We note that 14 and acyclicity are local conditions that can be checked and all solutions can be enumerated [^1] Recall that P-node skeletons have at most five edges and each R-node skeleton is a $K_{4}$. locally and independently for each skeleton. On the other hand 5 is a global property, which states that for each real edge $e$ that connects two P-nodes in a series, only one of them may put a crossing on $e$.
For each node $\mu$ of $\mathcal{T}$, let $\mathcal{F}_{\mu}$ be a subset of the embeddings of $\operatorname{skel}(\mu)$ that satisfy 14. We call $\mathcal{F}_{\mu}$ the feasible embeddings of $\operatorname{skel}(\mu)$. We are interested in whether we can choose for each node $\mu$ a feasible embedding $\mathcal{E}_{\mu}\in\mathcal{F}_{\mu}$ such that together they satisfy 5. We call such a choice of embeddings consistent.
To find a consistent choice of embeddings, we construct an auxiliary graph $H$. The vertex set of $H$ is $\bigcup_{\mu\in\mathcal{T}}\mathcal{F}_{\mu}$. We turn each set $\mathcal{F}_{\mu}$ into a clique and we connect two embedding choices of different P-nodes if choosing both of them simultaneously violates 5. It is then clear that $H$ contains an independent set whose size equals the number of nodes of $\mathcal{T}$ if and only if there exists a consistent choice of embeddings. In the following, we show that this can be decided in linear time by proving that $H$ has bounded treewidth (Cygan, Fomin, Kowalik, Lokshtanov, Marx, Pilipczuk, Pilipczuk, and Saurabh, 2015).
Lemma 5.
The auxiliary graph $H$ has bounded treewidth.
Proof.
We first consider the following construction. Let $T$ be a tree, let $I$ be an independent set of vertices in $T$, and let $c$ be a constant. Let $T^{\prime}$ be the graph obtained by connecting for each vertex in $I$ its neighbors by a cycle that visits them in some arbitrary order. We say that $T^{\prime}$ is a closure of $T$ with respect to $I$. Since $I$ is an independent set, each block of $T^{\prime}$ is either an edge or a wheel, and hence $T^{\prime}$ has treewidth at most $3$. Let $T^{\prime\prime}$ be obtained from $T^{\prime}$ by expanding each vertex $v$ of $T^{\prime}$ into a clique $C_{v}$ of size at most $c$ such that each edge $uv$ of $T^{\prime}$ is expanded into a biclique completely connecting $C_{u}$ and $C_{v}$. We call $T^{\prime\prime}$ the $c$-clique expansion of $T^{\prime}$. Clearly, the treewidth of $T^{\prime\prime}$ is at most $3c$.
We now use this to bound the treewidth of $H$. For each S-node $\lambda$ of $\mathcal{T}$ the skeleton $\operatorname{skel}(\lambda)$ defines a circular ordering of its virtual edges. Note that embeddings of two different P-nodes $\mu$ and $\mu^{\prime}$ can be connected by an edge only if they have a common S-node neighbor $\lambda$ such that their corresponding virtual edges in $\operatorname{skel}(\lambda)$ are consecutive in $\operatorname{skel}(\lambda)$. Therefore the auxiliary graph $H$ is a subgraph of the graph obtained from the SPQR-tree $\mathcal{T}$ by (i) forming the closure $T^{\prime}$ of $\mathcal{T}$ with respect to the (independent set of) S-nodes, where we connect the neighbors of the S-node in the order in which the corresponding virtual edges appear in the skeleton of the S-node and (ii) taking the $c$-clique-expansion of $T^{\prime}$ where $c$ is an upper bound on the number of embeddings satisfying C1–C4 for any skeleton. Auer et al. (2016) show that $c\leq 12$. Hence the treewidth of $H$ is at most $36$. ∎
Theorem 10.
There is a linear-time algorithm for testing whether a given single-source graph admits an upward outer-1-planar embedding.
Proof.
By Lemma 2 we may assume that $G$ is biconnected. We first compute the SPQR-tree $\mathcal{T}$ of $G$ in linear time (Gutwenger and Mutzel, 2000). Next, we check that the skeleton of each R-node of $\mathcal{T}$ is a $K_{4}$, and that the skeleton of each P-node of $\mathcal{T}$ has at most four virtual edges. If this fails, we can reject the instance as it does not have an outer-1-planar embedding (Auer et al., 2016). Next, we compute for each node $\mu$ of $\mathcal{T}$ the set $\mathcal{F}_{\mu}$ that contains all acyclic embeddings of $\operatorname{skel}(\mu)$ that satisfy conditions 14. Since skeletons of S-nodes have a unique planar embedding (2) and the skeletons of P- and R-nodes have bounded size, this can be done in total linear time. It then remains to consistently choose these embeddings so that also Condition 5 is satisfied. To this end, we construct the auxiliary graph $H$ and compute a maximum independent set, which takes linear time as well (Cygan et al., 2015). If the size of the maximum independent set is smaller than the number of nodes of $\mathcal{T}$, there is no consistent choice of embeddings and we reject the instance. Otherwise, by Theorem 9 this choice defines an outer-1-planar embedding $\mathcal{E}$. By Lemma 4 the embedding $\mathcal{E}$ has an acyclic planarization and is hence upward outer-1-planar by Theorem 8. We note that, in the positive case, the embedding $\mathcal{E}$ can also be constructed in linear time. ∎
7 Conclusion
In this paper we initiated the study of upward $k$-planar drawings, that is, upward drawings of directed acyclic graphs such that every edge is crossed at most $k$ times for a given constant $k$. We first gave upper and lower bounds for the upward local crossing number of various graph families, i.e., the minimum $k$ such that every graph from the respective family admits an upward $k$-planar drawing. We strengthen these combinatorial results by proving that testing a DAG for upward $k$-planarity is NP-complete even for $k=1$. On the positive side, testing upward outer-1-planarity for single source digraphs can be done in linear time. We conclude the paper by listing some open problems that may stimulate further research.
  1. 1.
    Is there a directed outerpath that does not admit an upward 1-planar drawing?
  2. 2.
    Consider the class $\mathcal{O}_{\Delta}$ of outerplanar graphs (or even 2-trees) of maximum degree $\Delta$. Is there a function $f$ such that every graph in $\mathcal{O}_{\Delta}$ admits an $f(\Delta)$-planar upward drawing?
  3. 3.
    In light of the lower bounds in Section 3, it is natural to consider graphs with a special structure, in order to prove sublinear upper bounds on their (upward) local crossing number. For example, Wood and Telle (Wood and Telle, 2007, Corollary 8.3) show that every (undirected) graph of maximum degree $\Delta$ and treewidth $\tau$ admits a (straight-line) drawing in which every edge crosses $\mathcal{O}(\Delta^{2}\tau)$ other edges. Can the upward local crossing number be bounded similarly by a function in $\Delta$ and $\tau$?
  4. 4.
    Do planar graphs of maximum degree $\Delta$ have upward local crossing number $\mathcal{O}(f(\Delta)n^{1-\epsilon})$ for some function $f$ and some constant $\epsilon>0$?
  5. 5.
    Can upward outer-1-planarity be efficiently tested for multi-source and multi-sink DAGs?
  6. 6.
    Investigate parameterized approaches to testing upward $1$-planarity.